OpenStudy (anonymous):

please help:) evaluate the following without using a calculator 1. (1^7/9)^3/2

OpenStudy (anonymous):

1^ any thing = 1, so first answer is 1 32^-4/5 means 1/(32)^4/5 which is 1/(fifth root of 32)^4 = 1/2^4 = 1/16

OpenStudy (anonymous):

i have the answers but i dont dont know the workings out and the first one isnt 1 its 2 10/27 :/ but the 2nd ones right :) thanks

OpenStudy (anonymous):

@gcuwagirl

OpenStudy (anonymous):

I misread the problem for 1 then. Is it 1 raised to the 7 and then divide that by 9 raised to the 3/2?

OpenStudy (anonymous):

yep